# Single-fluid simulation of partially-ionized, non-ideal plasma facilitated by a tabulated equation of state

### v1

@@ II.1 Single-fluid model of partially ionized plasma
with $\kappa$ as the thermal conductivity and $\nabla T$ as the temperature gradient. The fluids are coupled to one another through the source terms $\mathbf{R}$ and $Q$, which represent the collisional transfer of momentum and energy respectively.
@@
$\rho=\rho_{i}+\rho_{n},\quad\mathbf{v}=\frac{\rho_{i}v_{i}+\rho_{n}v_{n}}{\rho_{i}+\rho_{n}},\quad\zeta=\zeta_{ei}+\zeta_{n},\quad p=p_{ei}+p_{n},$
@@
$\zeta_{a}=\frac{1}{2}m_{a}v_{a}^{2}+\rho_{a}\varepsilon_{a}.$ (22)
@@ II.2 Simulated system
These modifications create a fully conservative system on the LHS, and all non-conservative, diffusive effects have been moved to the RHS to be treated as a source term through operator splitting, which allows our system to accurately capture discontinuities and shocks.
@@ IV Verification and demonstration
$w_{x}$, $w_{y}$ and $w_{z}$ are not included as they are always initialized to 0. The thermodynamic state is adjusted through the density and pressure scale factors $\alpha_{\rho}$ and $\alpha_{p}$. The rest of the variables are then rescaled accordingly through the rescaling vector $\mathbf{S}$:
$\mathbf{S}=(\alpha_{\rho},\frac{\sqrt{\alpha_{p}}}{\alpha_{\rho}},\frac{\sqrt{\alpha_{p}}}{\alpha_{\rho}},\frac{\sqrt{\alpha_{p}}}{\alpha_{\rho}},\alpha_{p},\sqrt{\alpha_{p}},\sqrt{\alpha_{p}},\sqrt{\alpha_{p}})^{T}.$
for the simulated quantities and the rescaling factor $\nicefrac{{\alpha_{\rho}}}{{\alpha_{p}}}$ for time. Altogether, the actual initialized state $\mathbf{u}^{*}_{0}$ and final time $t_{f}^{*}$ of the simulation are given as:
$\mathbf{u}^{*}_{0}=\mathbf{u}_{0}\cdot\mathbf{S},\quad\quad t_{f}^{*}=\frac{\alpha_{\rho}}{\alpha_{p}}t_{f}.$
@@ IV.2 Ion-neutral relative-motion
The initial state is kept constant for this test, with $\alpha_{\rho}=0.65\times 10^{-7}\text{kg}\cdot\text{m}^{-3}$, and $\alpha_{p}=10$ Pa, leading to an ionization fraction of $\sim$0.15. Instead, we vary the final time of the simulation to compare the degree of coupling between the ions and the neutrals for phenomena of different timescales. The $x$-axis has been normalized by the final time to facilitate this comparison.
@@
Our results agree with those of Hillier, Takasao, and Nakamura (2016), which were simulated using a two-fluid model (neutral and charged). We see that in the short timescale cases, in the region following the rarefaction, a significant velocity has only been induced in the ions, while $v_{n,x}$ and $v_{n,y}$ remain close to zero. However, since both species feature a common temperature, the slow shock also leads to the heating of the neutrals, and the creation of a steep pressure gradient in the $x$-direction. Since the neutrals in the pre-shock region are almost stationary, this highly localized heating leads to an explosive blast wave Needham (2010), characterized by a dramatic but localized jump in $v_{n,x}$. These results show that for a pressure of $p=10$ Pa, the charged and neutral dynamics are not equilibrated at the Alfvénic timescale, and hence cannot be simulated using only a MHD simulation, but can however be captured by our model.
@@ A.1 Specific heat ratio, $\gamma$
Fig. 7 shows the values of $\gamma$ mapped across the PIR. We see that as expected the variation of $\gamma$ very closely follows the variations in the ionization fraction. $\gamma$ reaches the monoatomic equipartition prediction of $\nicefrac{{5}}{{3}}$ at both the $\xi_{i}\rightarrow 1$ and $\xi_{i}\rightarrow 0$ limits. This is because there are minimal ionization events taking place in these limits, and the plasma resembles a monoatomic ideal gas chemically (our model does not current include diatomic particles). However, $\gamma$ changes rapidly with $\xi_{i}$ between the limits, reaching a minima of $\sim 1.15$. Hence, significant modifications of the plasma compressibility are expected in this region, necessitating the use of a more advanced EoS.
@@ A.2 Collisionality
Figure 8: Heat maps showing the variation of inter-species collision frequencies, $\nu_{ab}$, of a PIP with number density and temperature. The black contours are for $\nu$, while the white contours shows the ionization fraction, $\xi_{i}$, of the plasma. (A): ion-electron collision frequency, $\nu_{ei}$, (B): ion-neutral collision frequency, $\nu_{in}$, (C): electron-neutral collision frequency, $\nu_{en}$, (D): The ratio $\nicefrac{{\nu_{ei}}}{{\nu_{in}}}$, which indicates the relative collisional coupling strength between the charged particles and between the ions and the neutrals.
@@ A.4 Timescale comparisons
The mapping above can also be combined with the characteristics time and length scales of a specific partially ionized phenomena to reveal the dominant physics expected. Let us consider the example of Alfvénic timescale within tokamak reactors, which characterizes the growth rate of MHD instabilities. The Alfvén time $\tau_{A}$ is given as
@@
where $v_{A}$ is the Alfvén speed, and $\mu_{0}$ is the permeability of free space. The length scale $L$ of the phenomena is taken as the size of the fusion reactor. In particular, we will assume a configuration akin to the International Thermonuclear Experimental Reactor (ITER) , which has a magnetic field strength of $B=5$T, and a minor radius of $L=5$m Aymar, Barabaschi, and Shimomura (2002).
Figure 10: Heatmap showing variations in the ratio between the ion-neutral collision time, $\tau_{in}$, and the Alfvén time scale $\tau_{A}$, calculated with $B$ = 5T and length scale $L$ = 5m, with different thermodynamic states across the PIR.
Fig. 10 shows the comparisons between $\tau_{A}$ and $\tau_{in}$, which establishes whether the neutrals are coupled to the electromagnetically driven dynamics of the charged species. We see that in ITER-like conditions, a timescale of $\tau_{A}$ is insufficient for the collisional coupling of neutrals with the ions if $n\leq 10^{20}$. Both the tokamak core ($n\approx 10^{20}$) and edge ($n\approx 10^{19}$) are below this limit. Consequently, a simulation needs to capture the relative-motion between the ions and the neutrals to accurately simulate the growth of MHD instabilities involving neutrals in ITER.

### v2

@@ II.1 Single-fluid model of partially ionized plasma
with $\kappa$ as the thermal conductivity and $\nabla T$ as the temperature gradient. The subscripts $i,e,n$ define quantities for the ions, electrons and neutrals respectively. The fluids are coupled to one another through the source terms $\mathbf{R}$ and $Q$, which represent the collisional transfer of momentum and energy respectively.
@@
$\rho=\rho_{i}+\rho_{n},\quad\mathbf{v}=\frac{\rho_{i}\mathbf{v}_{i}+\rho_{n}\mathbf{v_{n}}}{\rho_{i}+\rho_{n}},\quad\zeta=\zeta_{ei}+\zeta_{n},\quad p=p_{ei}+p_{n},$
@@
$\zeta_{a}=\frac{1}{2}\rho_{a}v_{a}^{2}+\rho_{a}\varepsilon_{a}.$ (20)
@@ II.2 Simulated system
These modifications create a fully conservative system on the LHS for the fluid density, momentum and energy, without the need for exchange source terms between the different particle species due to ionization/recombination. Diffusive effects have been moved to the RHS to be treated as a source term through operator splitting, which allows our system to accurately capture discontinuities and shocks. We note that energy exchange with the external environment will still come into the system as source terms, such as the loss of energy due to ionization radiation for an optically thin plasma. This energy loss is not currently included in the model and will be addressed in future work.
@@ IV Verification and demonstration
$w_{x}$, $w_{y}$ and $w_{z}$ are not included as they are always initialized to 0. The thermodynamic state is adjusted through the density and pressure scale factors $\beta_{\rho}$ and $\beta_{p}$. The rest of the variables are then rescaled accordingly through the rescaling vector $\mathbf{S}$:
$\mathbf{S}=(\beta_{\rho},\frac{\sqrt{\beta_{p}}}{\beta_{\rho}},\frac{\sqrt{\beta_{p}}}{\beta_{\rho}},\frac{\sqrt{\beta_{p}}}{\beta_{\rho}},\beta_{p},\sqrt{\beta_{p}},\sqrt{\beta_{p}},\sqrt{\beta_{p}})^{T}.$
for the simulated quantities and the rescaling factor $\nicefrac{{\beta_{\rho}}}{{\beta_{p}}}$ for time. Altogether, the actual initialized state $\mathbf{u}^{*}_{0}$ and final time $t_{f}^{*}$ of the simulation are given as:
$\mathbf{u}^{*}_{0}=\mathbf{u}_{0}\cdot\mathbf{S},\quad\quad t_{f}^{*}=\frac{\beta_{\rho}}{\beta_{p}}t_{f}.$
@@ IV.2 Ion-neutral relative-motion
The initial state is kept constant for this test, with $\alpha_{\rho}=0.65\times 10^{-7}\text{kg}\cdot\text{m}^{-3}$, and $\beta_{p}=10$ Pa, leading to an ionization fraction of $\sim$0.15. Instead, we vary the final time of the simulation to compare the degree of coupling between the ions and the neutrals for phenomena of different timescales. The $x$-axis has been normalized by the final time to facilitate this comparison.
@@
Our results agree with the two-fluid simulations (neutral and charged) by Hillier et al.Hillier, Takasao, and Nakamura (2016). We see that in the short timescale cases, in the region following the rarefaction, a significant velocity has only been induced in the ions, while $v_{n,x}$ and $v_{n,y}$ remain close to zero. However, since both species feature a common temperature, the slow shock also leads to the heating of the neutrals, and the creation of a steep pressure gradient in the $x$-direction. Since the neutrals in the pre-shock region are almost stationary, this highly localized heating leads to an explosive blast wave Needham (2010), characterized by a dramatic but localized jump in $v_{n,x}$. These results show that for a pressure of $p=10$ Pa, the charged and neutral dynamics are not equilibrated at the Alfvénic timescale, and hence cannot be simulated using only a MHD simulation, but can however be captured by our model.
@@ A.1 Specific heat ratio, $\gamma$
Fig. 7 shows the values of $\gamma$ mapped the partially-ionized regime. We see that as expected the variation of $\gamma$ very closely follows the variations in the ionization fraction. $\gamma$ reaches the monoatomic equipartition prediction of $\nicefrac{{5}}{{3}}$ at both the $\xi_{i}\rightarrow 1$ and $\xi_{i}\rightarrow 0$ limits. This is because there are minimal ionization events taking place in these limits, and the plasma resembles a monoatomic ideal gas chemically (our model does not current include diatomic particles). However, $\gamma$ changes rapidly with $\xi_{i}$ between the limits, reaching a minima of $\sim 1.15$. Hence, significant modifications of the plasma compressibility are expected in this region, necessitating the use of a more advanced EoS.
@@ A.2 Collisionality
Figure 8: Heat maps showing the variation of inter-species collision frequencies, $\nu_{ab}$, of a PIP with number density and temperature. The black contours are for $\nu$, while the white contours shows the ionization fraction, $\xi_{i}$, of the plasma. (A): ion-electron collision frequency, $\nu_{ei}$, (B): ion-neutral collision frequency, $\nu_{in}$, (C): electron-neutral collision frequency, $\nu_{en}$, (D): The ratio $\frac{m_{e}\nu_{ei}}{m_{i}\nu_{in}}$, which indicates the relative collisional coupling strength between the charged particles and between the ions and the neutrals.
@@ A.4 Timescale comparisons
The analysis described above can also be combined with the characteristics time and length scales of a specific partially-ionized phenomena to reveal the dominant physics expected. Let us consider the example of Alfvénic timescale within tokamak reactors, which characterizes the growth rate of MHD instabilities. The Alfvén time $\tau_{A}$ is given as
@@
where $v_{A}$ is the Alfvén speed, and $\mu_{0}$ is the permeability of free space. The length scale $L$ of the phenomena is taken as the size of the fusion reactor. In particular, we will assume a configuration akin to the International Thermonuclear Experimental Reactor (ITER) , which has a magnetic field strength of $B\approx 5$T, and a minor radius of $L\approx 3$m Aymar, Barabaschi, and Shimomura (2002).
Figure 10: Heatmap showing variations in the ratio between the ion-neutral collision time, $\tau_{in}$, and the Alfvén time scale $\tau_{A}$, calculated with an ITER-like configuration of $B$ = 5T and length scale $L$ = 5m, with different thermodynamic states across the PIR.
Fig. 10 shows the comparisons between $\tau_{A}$ and $\tau_{in}$, which establishes whether the neutrals are coupled to the electromagnetically driven dynamics of the charged species. We see that in ITER-like conditions, a timescale of $\tau_{A}$ is insufficient for the collisional coupling of neutrals with the ions if $n\leq 10^{20}$. Both the tokamak core ($n\approx 10^{20}$) and edge ($n\approx 10^{19}$) are below this limit. Consequently, a simulation needs to account for the relative-motion between the ions and the neutrals to accurately simulate the growth of MHD instabilities involving neutrals, justifying the need for our single-fluid model over traditional extended-MHD models.